OpenStudy (anonymous):

6x^2+5=7x i have to use quadratic formula to find exact solution?

OpenStudy (bahrom7893):

first rewrite: 6x^2 -7x +5=0 now this is in the form: ax^2 + bx +c = 0 where a = 6, b = -7 and c = 5 Now use the formula: x1 = [ -b-sqrt(b^2-4ac) ]/ [2a] x2 = [ -b+sqrt(b^2-4ac) ]/ [2a]

OpenStudy (anonymous):

i don't understand the x1 = [ -b-sqrt(b^2-4ac) ]/ [2a] x2 = [ -b+sqrt(b^2-4ac) ]/ [2a] part :S

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

x = -7 +- sqrt(-7^2-4*6*5)/(2*6) x= -7 +- sqrt(-71)/12 x= -7 +- i sqrt(71)/12